OpenStudy (anonymous):

12.) Refer to right triangle ACB. If m Ang A = 30° and BC = 5, what is the length of AB? 5 sqrt(2) 5 5 sqrt(3) 10

OpenStudy (anonymous):

A |\ | \ | \ |__\ C B sin(30) = 5/ x x = 5 / sin(30) x = 2.5

OpenStudy (anonymous):

2.5 doesn't match with any of the answers but thank you anyways.

OpenStudy (anonymous):

A |\ | \ | \ |__\ B C tan(30) = 5/x x = 5 / tan(30) x = 5/ sqrt 3 still not the answer... T-T